OpenStudy (bloomlocke367):

Hmmm...this is a dumb question, but to get rid of a fraction in an equation, do you multiply everything by the denominator of the fraction, or the reciprocal?

OpenStudy (bloomlocke367):

@jdoe0001 lol sorry for bothering you

OpenStudy (anonymous):

no question is a dumb question

Nnesha (nnesha):

you would multiply `with the reciprocal` of the denominator

OpenStudy (bloomlocke367):

It is when you're in precal honors and can't remember basic math rules lol XD

OpenStudy (bloomlocke367):

thank you

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

im honers to

Nnesha (nnesha):

nice!

OpenStudy (anonymous):

:3

OpenStudy (jdoe0001):

there was a fellow the other day here he said he knew physics, but never seen a logarithm :|

OpenStudy (jdoe0001):

go figure

OpenStudy (anonymous):

r u Muslim Nnesha

OpenStudy (anonymous):

i am

OpenStudy (bloomlocke367):

So I have \(x^2-7x-\frac{3}{4}=0\), and to get rid of that fraction I'd multiply everything} by \(\frac{4}{3}\)?

Nnesha (nnesha):

ahh no!

OpenStudy (bloomlocke367):

wait...you said reciprocal of the denominator... so 1/4

OpenStudy (bloomlocke367):

yes, no, maybe so?

OpenStudy (bloomlocke367):

Oh, and @jdoe0001, I'm taking AP Physics right now XD I love it

Nnesha (nnesha):

yes right

OpenStudy (bloomlocke367):

That just creates more fractions D:

Nnesha (nnesha):

:3 x^2-7x-\frac{ 3 }{ 4} \times \frac{1}{4} \[x^2-7x-\frac{ 3 }{ 4} \times \frac{1}{4} \] don't see any points to get more messy stuff :P

OpenStudy (triciaal):

there are no dumb questions!!!!

Nnesha (nnesha):

i didn't noticed it's 1/4 not 4/3

OpenStudy (anonymous):

thats what i sayed

OpenStudy (bloomlocke367):

I dunno...

Nnesha (nnesha):

ignore me plz :(

OpenStudy (bloomlocke367):

ummm... okay. I guess I'll just ask Mrs. Dunn tomorrow

Nnesha (nnesha):

no wait don't close !

OpenStudy (anonymous):

nesha r u muslim i am

OpenStudy (bloomlocke367):

Please don't put spam on my question. >.<

OpenStudy (jdoe0001):

\(\bf x^2-7x-\frac{3}{4}=0\implies 4\left( x^2-7x-\frac{3}{4} \right)=4(0) \\ \quad \\ 4x^2-21x-\cancel{4}\cdot \cfrac{3}{\cancel{4}}=0\implies 4x^2-21x-3=0\)

OpenStudy (triciaal):

to get rid of a fraction in an equation, do you multiply everything by the denominator of the fraction, or the reciprocal? you multiply everything by the lowest common denominator

OpenStudy (jdoe0001):

hmmm rather \(\bf x^2-7x-\frac{3}{4}=0\implies 4\left( x^2-7x-\frac{3}{4} \right)=4(0) \\ \quad \\ 4x^2-28x-\cancel{4}\cdot \cfrac{3}{\cancel{4}}=0\implies 4x^2-28x-3=0\)

OpenStudy (bloomlocke367):

OKAY THANK YOU OMG.

OpenStudy (triciaal):

remember the denominator of a whole number is 1

OpenStudy (triciaal):

you are welcome

OpenStudy (bloomlocke367):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @triciaal remember the denominator of a whole number is 1 \(\color{#0cbb34}{\text{End of Quote}}\) Yes, I know :D

Nnesha (nnesha):

fraction in the equation i was doing totally different thin'!

Nnesha (nnesha):

srry about that

OpenStudy (bloomlocke367):

It's okay :) thanks guys